OpenStudy (anonymous):

Find the vertex of y = 2(x+3)(x-1) ? explain how to do it

OpenStudy (anonymous):

calculus or geometry? the answer depends

OpenStudy (anonymous):

im in algebra 2

OpenStudy (anonymous):

There are multiple ways to do it; This is because a parabola is horizontally symmetric alone its center. First way find x-intercepts: x = -3 and x = 1 now since its symmetric you know the x value of the vertex is equidistant to those 2 values; therefore x(vertex) = (-3+1)/2 = -1 y(vertex) = 2*2*-2 = -8 You can also find it other ways but I suppose you don't need them here.

OpenStudy (dan815):